Doreen Schmidt is a chemist. She needs to prepare 28 ounces of a 10​% hydrochloric acid solution. Find the amount of 14​% solution and the amount of 7​% solution she should mix to get this solution.
Answer by ankor@dixie-net.com(22740)

prepare 28 ounces of a 10​% hydrochloric acid solution.
Find the amount of 14​% solution and the amount of 7​% solution she should mix to get this solution.
:
let x = amt of 14% solution required
the total amt should be 28 oz, therefore:
(28-x) = amt of 7% solution
:
A typical mixture equation
.14x + .07(28-x) = .10(28)
.14x + 1.96 - .07x = 2.8
.14x - .07x = 2.8 - 1.96
.07x = .84
x = .84/.07
x = 12 oz of 14% solution
then
28 - 12 = 16 oz of the 7% solution
:
:
Check this
.14(12) + .07(16) = .10(28)
1.68 + 1.12 = 2.8